# MME-RealWorld: Could Your Multimodal LLM Challenge High-Resolution Real-World Scenarios that are Difficult for Humans?

### v1

@@ Abstract
Comprehensive evaluation of Multimodal Large Language Models (MLLMs) has recently garnered widespread attention in the research community. However, we observe that existing benchmarks present several common barriers that make it difficult to measure the significant challenges that models face in the real world, including: 1) small data scale leads to a large performance variance; 2) reliance on model-based annotations results in restricted data quality; 3) insufficient task difficulty, especially caused by the limited image resolution. To tackle these issues, we introduce MME-RealWorld. Specifically, we collect more than $300$K images from public datasets and the Internet, filtering $13,366$ high-quality images for annotation. This involves the efforts of professional $25$ annotators and $7$ experts in MLLMs, contributing to $29,429$ question-answer pairs that cover $43$ subtasks across $5$ real-world scenarios, extremely challenging even for humans. As far as we know, MME-RealWorld is the largest manually annotated benchmark to date, featuring the highest resolution and a targeted focus on real-world applications. We further conduct a thorough evaluation involving $28$ prominent MLLMs, such as GPT-4o, Gemini 1.5 Pro, and Claude 3.5 Sonnet. Our results show that even the most advanced models struggle with our benchmarks, where none of them reach 60% accuracy. The challenges of perceiving high-resolution images and understanding complex real-world scenarios remain urgent issues to be addressed. The data and evaluation code are released in our Project Page.
@@ 1 Introduction
In light of these concerns, we propose a new benchmark named MME-RealWorld. We first pay attention to a series of well-motivated families of datasets, considering images from sources such as autonomous driving, remote sensing, video surveillance, newspapers, street views, and financial charts. These scenarios are difficult even for humans, where we hope that MLLMs can really help. Considering these topics, we collect a total of $13,366$ high-resolution images from more than $300$K public and internet sources. These images have an average resolution of $2,000$$\times$$1,500$, containing rich image details. $25$ professional annotators and $7$ experts in MLLMs are participated to annotate and check the data quality, and meanwhile ensuring that all questions are challenging for MLLMs. Note that most questions are even hard for humans, requiring multiple annotators to answer and double-check the results. As shown in Fig. 2(a), MME-RealWorld finally contains $29,429$ annotations for $43$ sub-class tasks, where each one has at least $100$ questions. $28$ advanced MLLMs are evaluated on our benchmark, along with detailed analysis. We conclude the main advantages of MME-RealWorld compared to existing counterparts as follows:
@@
2 Related Work
Multimodal Benchmark. With the development of MLLMs, a number of benchmarks have been built. For instance, MME [17] constructs a comprehensive evaluation benchmark that includes a total of 14 perception and cognition tasks. All QA pairs in MME are manually designed to avoid data leakage, and the binary choice format makes it easy to quantify. MMBench [45] contains over $3,000$ multiple-choice questions covering $20$ different ability dimensions, such as object localization and social reasoning. It introduces GPT-4-based choice matching to address the MLLM’s lack of instruction-following capability and a novel circular evaluation strategy to improve the evaluation robustness. Seed-Bench [34] is similar to MME and MMBench but consists of $19,000$ multiple-choice questions. The larger sample size allows it to cover more ability aspects and achieve more robust results. SEED-Bench-2 [33] expands the dataset size to $24,371$ QA pairs, encompassing $27$ evaluation dimensions and further supporting the evaluation of image generation. MMT-Bench [69] scales up the dataset even further, including $31,325$ QA pairs from various scenarios such as autonomous driving and embodied AI. It encompasses evaluations of model capabilities such as visual recognition, localization, reasoning, and planning. Additionally, other benchmarks focus on real-world usage scenarios [20, 48, 6] and reasoning capabilities [70, 5, 23]. However, there are widespread issues, such as data scale, annotation quality, and task difficulty, in these benchmarks, making it hard to assess the challenges that MLLMs face in the real world.
MLLMs. This field has undergone significant evolution [68, 19], initially rooted in BERT-based language decoders and later incorporating advancements in LLMs. MLLMs exhibit enhanced capabilities and performance, particularly through end-to-end training techniques, by leveraging advanced LLMs such as GPTs [53, 7], LLaMA [62, 63], Alpaca [59], PaLM [14, 2], BLOOM [52], Mistral [28], and Vicuna [13]. Recent model developments, including Flamingo [3], PaLI [29], PaLM-E [16], BLIP-2 [36], InstructBLIP [15], Otter [31], MiniGPT-4 [75], mPLUG-Owl [67], LLaVA [43], Qwen-VL [4], and VITA [18], bring unique perspectives to challenges such as scaling pre-training, enhancing instruction-following capabilities, and overcoming alignment issues. However, the performance of these models in the face of real scenarios has often not been revealed.
High-resolution MLLMs. Empirical studies have shown that employing higher resolution is an effective solution for many tasks [4, 41, 40, 51]. Approaches like LLaVA-Next [42] segment high-resolution images into multiple patches, encoding each one independently before concatenating all local patch tokens with the original global image tokens, albeit at an escalated computational cost. Other models, such as Monkey [40] and LLaVA-UHD [64], also split images into patches but subsequently compress them to avoid redundant tokens. Mini-Genimi [39] comprises twin encoders, one for high-resolution images and the other for low-resolution visual embedding. They work in an attention mechanism, where the low-resolution encoder generates visual queries, and the high-resolution counterpart provides candidate keys and values for reference. Conv-LlaVA [22] employs ConvNeXt instead of ViT as the vision encoder. Cambrian [61] uses a set of learnable latent queries that interact with multiple vision features via cross-attention layers. Additionally, SliME [73] stresses the importance of global features, compressing the local image patches twice but preserving all the global context. Although many of these models focus on improving resolution, none have been tested in a rigorous high-resolution benchmark, often providing only intuitive examples that lack informativeness and convincing results. Our proposed benchmark offers a rigorous evaluation to test the capabilities in understanding high-resolution images.
@@ 4 Experiments
We evaluate a total of 24 open-source MLLMs, including Qwen-VL-Chat [4], LLaVA, LLaVA-Next [30], TextMonkey [46], mPLUG-DocOwl 1.5 [25], ShareGPT4V [11], MiniGPT-v2 [9], Monkey [40], OtterHD [31], Cambrian-1 [61], Mini-Gemini-HD [39], MiniCPM-V 2.5 [26], DeepSeek-VL [47], YI-VL-34B [^1] https://huggingface.co/01-ai/Yi-VL-34B, SliME [73], CogVLM2222https://github.com/THUDM/CogVLM2, InternLM-XComposer2.5 [72], InternVL-Chat V1-5, and InternVL-2 [12], as well as 4 close-source MLLMs, including, GPT-4o [^3] https://openai.com/index/hello-gpt-4o/, GPT-4o-mini, Gemini 1.5 pro [60], and Claude 3.5 Sonnet [^4] https://www.anthropic.com/news/claude-3-5-sonnet.
@@ 4.1.1 Perception
1. GPT-4o performs best in real-world OCR tasks, achieving $77\%$ accuracy, but its performance drops significantly in more challenging tasks, falling behind other top-ranked models. This trend is also observed in other closed-source models, such as Gemini-1.5-Pro and GPT-4o-mini, which perform well in OCR tasks but struggle significantly in other real-world tasks. There are three possible reasons: 1) Close-source models often have limitations on the maximum image size and resolution when uploading local images. For example, Claude 3.5 Sonnet has a maximum resolution limit of $8$K and a maximum image quality of $5$MB, while GPT-4o and Gemini-pro allow up to $20$MB. This restricts the input of some high-quality images, as we have to compress the images for upload. 2) Close-source models tend to be more conservative. We observe that the proportion of responses, where closed-source models output “E” indicating that the object in question is not present in the image, is high. This suggests that these models may adopt a conservative response strategy to avoid hallucinations or to provide safer answers. 3) Closed-source models sometimes refuse to answer certain questions. Due to different input/output filtering strategies, some samples are considered to involve privacy or harmful content and are therefore not answered.
2. Models allowing higher resolution input, such as Mini-Gemini-HD and SliME, demonstrate a significant advantage over models directly using vision encoders like CLIP, such as ShareGPT4V and LLaVA1.5. At the same model size, these models consistently improve across different subtasks. This highlights the critical importance of high-resolution image processing for addressing complex real-world tasks.
3. There are also notable trends across different domains. Remote sensing tasks involve processing extremely large images, demanding a deeper comprehension of image details. Models that focus on high-resolution input, such as Cambrian-1, Mini-Gemini-HD, and SliME, outperform other models in these tasks. Additionally, models trained on large amounts of chart data exhibit improved perception capabilities for complex charts. For instance, SliME and LLaVA1.5 have limited and relatively simple chart data in their training sets, resulting in inferior performance in this category compared to more recent models.
@@ 4.1.2 Reasoning
Experimental results on the reasoning tasks are shown in Tab. 4. In terms of reasoning ability, Claude 3.5 Sonnet distinguishes itself as the top performer across most domains, particularly outpacing the second-place GPT-4o by $16.4\%$ in chart-related tasks. The closed-source model GPT-4o also performs well, trailing slightly behind the second-place InternVL-2 but even outperforming it in several domains. Most open-source models perform poorly, with traditional baseline methods such as LLaVA1.5 and Qwen-VL-Chat yielding results close to random guessing. Furthermore, reasoning tasks are more challenging than perception tasks. Even the top-ranked model fails to achieve an average accuracy above $45\%$, with class-based accuracy not exceeding $50\%$. This indicates that current models still have a significant gap to bridge to reach human-level reasoning capabilities.
@@ 4.2 Results on MME-RealWorld-CN
InternVL-2 InternLM2.5-7B-Chat 69.92 41.33 71.63 53.19 34.14 59.70 54.04
@@
1) InternVL-2 significantly outperforms existing models in both perception and reasoning tasks in the Chinese version, even surpassing its performance on the English version, indicating that it has been specifically optimized for Chinese data.
@@
InternVL-2 InternLM2.5-7B-Chat 61.90 61.20 41.79 31.92 44.12 49.20
Claude 3.5 Sonnet - 57.40 39.00 43.57 29.84 38.74 42.45
SliME-8B Llama3-8B 61.40 44.80 36.51 26.41 37.61 42.28
InternVL-Chat-V1-5 InternLM2-Chat-20B 54.00 32.80 41.16 31.18 37.25 39.79
CogVLM2-llama3-Chat Llama3-8B 56.80 35.40 37.35 28.94 36.48 39.62
YI-VL-34B Yi-34B-Chat 53.20 27.40 42.37 30.73 36.16 38.43
Monkey Qwen-7B 53.20 29.40 36.14 31.55 35.80 37.57
Mini-Gemini-7B-HD Vicuna-7B-v1.5 44.00 31.80 36.95 31.03 34.50 35.95
Mini-Gemini-34B-HD Nous-Hermes-2-Yi-34B 41.00 39.00 33.13 30.80 34.46 35.98
LLaVA-Next Llama3-8B 53.40 41.00 17.67 29.99 33.90 35.52
Cambrian-1-8B Llama3-8B-Instruct 47.00 39.80 25.81 26.79 32.48 34.85
SliME-13B Vicuna-13B 42.40 26.00 31.33 31.55 32.45 32.82
LLaVA-Next Qwen-72B 55.20 23.40 21.08 30.73 32.06 32.60
InternLM-XComposer2.5 InternLM2-7B 59.20 39.20 20.48 22.84 31.73 35.43
GPT-4o - 52.70 33.20 28.33 19.20 29.19 33.36
DeepSeek-VL DeepSeek-LLM-7b-base 27.20 20.80 27.31 33.04 28.84 27.09
Cambrian-1-34B Nous-Hermes-2-Yi-34B 45.20 23.80 16.67 27.31 27.98 28.25
LLaVA1.5-13B Vicuna-13B 17.20 34.20 27.31 29.69 27.86 27.10
ShareGPT4V-13B Vicuna-13B 55.00 36.00 19.48 16.07 27.06 31.64
MiniCPM-V 2.5 Llama3-8B 42.60 19.80 20.48 26.04 26.88 27.23
LLaVA1.5-7B Vicuna-7B 35.40 24.60 25.90 23.29 26.12 27.30
ShareGPT4V-7B Vicuna-7B 30.20 20.80 27.51 24.78 25.51 25.82
Qwen-VL-Chat Qwen-7B 26.00 20.80 27.31 24.55 24.63 24.67
GPT-4o-mini - 26.00 20.60 25.90 24.18 24.17 24.17
MiniGPT-v2 Llama 2-7B-Chat 24.15 20.60 26.10 24.18 23.88 23.76
mPLUG-DocOwl 1.5 LLama-7B 30.00 20.40 16.87 23.66 23.01 22.73
TextMonkey Qwen-7B 28.60 13.60 16.47 24.63 21.95 20.83
Gemini-1.5-pro - 30.40 2.20 4.42 20.01 15.96 14.26
@@ 4.3 Fine-grained Analysis and Findings
Existing Models Still Lacking in Image Detail Perception. Fig. 3 displays the frequency with which various models choose “E” as their answer. We compare $4$ close-source models with the best-performing open-source model, InternVL-2. During our annotation process, the frequency of “E” answers does not exceed $5\%$ of the overall data, meaning it represents only a small portion of the total QA pairs. However, nearly all models show a much higher frequency of “E” outputs than the actual number of “E” instances present in our benchmark. This indicates that most models’ visual perception modules fail to identify the objects in the images corresponding to our questions.

### v3

@@ Abstract
Comprehensive evaluation of Multimodal Large Language Models (MLLMs) has recently garnered widespread attention in the research community. However, we observe that existing benchmarks present several common barriers that make it difficult to measure the significant challenges that models face in the real world, including: 1) small data scale leads to a large performance variance; 2) reliance on model-based annotations results in restricted data quality; 3) insufficient task difficulty, especially caused by the limited image resolution. To tackle these issues, we introduce MME-RealWorld. Specifically, we collect more than $300$ K images from public datasets and the Internet, filtering $13,366$ high-quality images for annotation. This involves the efforts of professional $25$ annotators and $7$ experts in MLLMs, contributing to $29,429$ question-answer pairs that cover $43$ subtasks across $5$ real-world scenarios, extremely challenging even for humans. As far as we know, MME-RealWorld is the largest manually annotated benchmark to date, featuring the highest resolution and a targeted focus on real-world applications. We further conduct a thorough evaluation involving $29$ prominent MLLMs, such as GPT-4o, Gemini 1.5 Pro, and Claude 3.5 Sonnet. Our results show that even the most advanced models struggle with our benchmarks, where none of them reach 60% accuracy. The challenges of perceiving high-resolution images and understanding complex real-world scenarios remain urgent issues to be addressed. The data and evaluation code are released in our Project Page.
@@ 1 Introduction
In light of these concerns, we propose a new benchmark named MME-RealWorld. We first pay attention to a series of well-motivated families of datasets, considering images from sources such as autonomous driving, remote sensing, video surveillance, newspapers, street views, and financial charts. These scenarios are difficult even for humans, where we hope that MLLMs can really help. Considering these topics, we collect a total of $13,366$ high-resolution images from more than $300$K public and internet sources. These images have an average resolution of $2,000$$\times$$1,500$, containing rich image details. $25$ professional annotators and $7$ experts in MLLMs are participated to annotate and check the data quality, and meanwhile ensuring that all questions are challenging for MLLMs. Note that most questions are even hard for humans, requiring multiple annotators to answer and double-check the results. As shown in Fig. 2(a), MME-RealWorld finally contains $29,429$ annotations for $43$ sub-class tasks, where each one has at least $100$ questions. $29$ advanced MLLMs are evaluated on our benchmark, along with detailed analysis. We conclude the main advantages of MME-RealWorld compared to existing counterparts as follows:
@@ 3 Experiments
We evaluate a total of 24 open-source MLLMs, including Qwen-VL-Chat [4], LLaVA, LLaVA-Next [30], TextMonkey [46], mPLUG-DocOwl 1.5 [25], ShareGPT4V [11], MiniGPT-v2 [9], Monkey [40], OtterHD [31], Cambrian-1 [61], Mini-Gemini-HD [39], MiniCPM-V 2.5 [26], DeepSeek-VL [47], YI-VL-34B [^1] https://huggingface.co/01-ai/Yi-VL-34B, SliME [73], CogVLM2222https://github.com/THUDM/CogVLM2, InternLM-XComposer2.5 [72], InternVL-Chat V1-5, InternVL-2 [12], and Qwen2-VL [^3] https://github.com/QwenLM/Qwen2-VL, as well as 4 close-source MLLMs, including, GPT-4o [^4] https://openai.com/index/hello-gpt-4o/, GPT-4o-mini, Gemini 1.5 pro [60], and Claude 3.5 Sonnet [^5] https://www.anthropic.com/news/claude-3-5-sonnet.
@@ 3.1.1 Perception
  • GPT-4o performs well in real-world OCR tasks, achieving 77% accuracy, which is second only to Qwen2-VL. However, its performance significantly drops in more challenging tasks, lagging behind other top-ranked models. This trend is also observed in other closed-source models, such as Gemini-1.5-Pro and GPT-4o-mini, which perform well in OCR tasks but struggle significantly in other real-world tasks. There are three possible reasons: 1) Close-source models often have limitations on the maximum image size and resolution when uploading local images. For example, Claude 3.5 Sonnet has a maximum resolution limit of $8$K and a maximum image quality of $5$MB, while GPT-4o and Gemini-pro allow up to $20$MB. This restricts the input of some high-quality images, as we have to compress the images for upload. 2) Close-source models tend to be more conservative. We observe that the proportion of responses, where closed-source models output “E” indicating that the object in question is not present in the image, is high. This suggests that these models may adopt a conservative response strategy to avoid hallucinations or to provide safer answers. 3) Closed-source models sometimes refuse to answer certain questions. Due to different input/output filtering strategies, some samples are considered to involve privacy or harmful content and are therefore not answered.
  • Models allowing higher resolution input, such as Mini-Gemini-HD and SliME, demonstrate a significant advantage over models directly using vision encoders like CLIP, such as ShareGPT4V and LLaVA1.5. At the same model size, these models consistently improve across different subtasks. This highlights the critical importance of high-resolution image processing for addressing complex real-world tasks.
  • There are also notable trends across different domains. Remote sensing tasks involve processing extremely large images, demanding a deeper comprehension of image details. Models that focus on high-resolution input, such as Cambrian-1, Mini-Gemini-HD, and SliME, outperform other models in these tasks. Additionally, models trained on large amounts of chart data exhibit improved perception capabilities for complex charts. For instance, SliME and LLaVA1.5 have limited and relatively simple chart data in their training sets, resulting in inferior performance in this category compared to more recent models.
@@ 3.1.2 Reasoning
Experimental results on the reasoning tasks are shown in Tab. 4. In terms of reasoning ability, Claude 3.5 Sonnet distinguishes itself as the top performer across most domains, particularly outpacing the second-place Qwen2-VL by $12.6\%$ in chart-related tasks. The closed-source model GPT-4o also performs well, trailing slightly behind the third-place InternVL-2 but even outperforming it in several domains. Most open-source models perform poorly, with traditional baseline methods such as LLaVA1.5 and Qwen-VL-Chat yielding results close to random guessing. Furthermore, reasoning tasks are more challenging than perception tasks. Even the top-ranked model fails to achieve an average accuracy above $45\%$, with class-based accuracy not exceeding $50\%$. This indicates that current models still have a significant gap to bridge to reach human-level reasoning capabilities.
@@ 3.2 Results on MME-RealWorld-CN
Qwen2-VL Qwen-7B 70.28 38.33 89.20 29.40 36.86 59.80 52.81
InternVL-2 InternLM2.5-7B-Chat 69.92 41.33 71.63 39.3 34.14 57.97 51.26
@@
1) Qwen2-VL and InternVL-2 significantly outperform existing models in both perception and reasoning tasks in the Chinese version. The performance of these two models even surpasses their performance on the English version of MME-RealWorld, indicating that they jave been specifically optimized for Chinese data.
@@
InternVL-2 InternLM2.5-7B-Chat 44.93 74.92 38.14 29.00 46.65 46.75
Qwen2-VL Qwen2-7B 38.16 72.92 57.00 33.37 46.46 50.36
Claude 3.5 Sonnet - 74.44 65.79 31.54 25.12 44.31 49.22
SliME-8B LLama3-8B 44.44 70.93 30.54 29.13 44.21 43.76
InternVL-Chat-V1-5 InternLM2-Chat-20B 48.79 67.11 30.20 29.88 43.74 44.00
CogVLM2-llama3-Chat LLama3-8B 33.81 65.24 37.25 29.00 42.25 41.33
YI-VL-34B Yi-34B-Chat 37.68 61.46 33.22 29.75 41.16 40.53
Monkey Qwen-7B 43.96 56.81 32.55 28.38 39.70 40.43
Mini-Gemini-7B-HD Vicuna-7B-v1.5 28.50 68.61 25.50 24.00 38.80 36.65
Mini-Gemini-34B-HD Nous-Hermes-2-Yi-34B 29.95 60.47 25.50 29.63 38.75 36.39
LLaVA-Next LLama3-8B 14.93 58.14 29.53 28.25 36.44 32.71
Cambrian-1-8B LLama3-8B-Instruct 27.54 47.51 31.21 31.25 35.97 34.38
SliME-13B Vicuna-13B 44.93 49.17 30.54 23.87 35.18 37.13
LLaVA-Next Qwen-72B 24.64 40.87 27.52 28.12 31.67 30.29
InternLM-XComposer2.5 InternLM2-7B 18.36 40.70 16.78 30.00 30.05 26.46
GPT-4o - 33.81 39.87 20.81 22.75 29.05 29.31
DeepSeek-VL DeepSeek-LLM-7b-base 36.23 23.59 25.50 29.25 27.63 28.64
Cambrian-1-34B Hermes2-Yi-34B 21.74 31.40 23.49 25.12 26.48 25.44
LLaVA1.5-13B Vicuna-13B 36.23 27.08 25.84 23.25 26.27 28.10
ShareGPT4V-13B Vicuna-13B 35.75 27.91 24.83 22.88 26.17 27.84
MiniCPM-V 2.5 LLama3-8B 36.23 29.90 16.44 23.87 25.95 26.61
LLaVA1.5-7B Vicuna-7B 33.33 25.91 25.17 23.25 25.48 26.92
ShareGPT4V-7B Vicuna-7B 33.33 25.91 24.83 23.25 25.43 26.83
Qwen-VL-Chat Qwen-7B 30.92 36.41 13.42 19.88 25.29 25.16
GPT-4o-mini - 27.88 27.08 14.77 26.87 25.16 24.15
MiniGPT-v2 Llama 2-7B-Chat 34.30 28.57 19.80 22.13 25.12 26.20
mPLUG-DocOwl 1.5 LLama-7B 37.68 24.42 19.80 22.38 24.28 26.07
TextMonkey Qwen-7B 27.53 31.07 12.08 22.50 24.12 23.29
Gemini-1.5-pro - 5.30 5.32 14.77 15.67 11.14 10.26
@@ 3.3 Fine-grained Analysis and Findings
Existing Models Still Lacking in Image Detail Perception. Fig. 3 displays the frequency with which various models choose “E” as their answer. We compare $4$ close-source models with the top-level open-source model, InternVL-2. During our annotation process, the frequency of “E” answers does not exceed $5\%$ of the overall data, meaning it represents only a small portion of the total QA pairs. However, nearly all models show a much higher frequency of “E” outputs than the actual number of “E” instances present in our benchmark. This indicates that most models’ visual perception modules fail to identify the objects in the images corresponding to our questions.